Direction: Answer the following and illustrate each under the normal curve:


1. Compute the probability area to the left of z = -1.25.


2. Compute the probability area above z = 1.


3. Find the probability area between z = -0.25 and z = 1.5.


4. Find the 90th percentile of a normal curve.


5. Compute the upper 5% of the normal curve.

1
Expert's answer
2022-03-28T17:04:36-0400

1) "P(Z<-1.25)=0.10565"

2) "P(Z>1)=0.15866"




3) "P(-0.25<Z<1.5)=P(Z<1.5)-P(Z<-0.25)=0.93319-0.40129=0.5319"




4) "P(Z<a)=0.9\implies a=1.282"






5) The upper 5% is interval "(a,+\infty)" where

"P(Z>a)=0.05\implies a=1.645"